# Closing the stellar labels gap: Stellar label independent evidence for [/M] information in Gaia BP/RP spectra

### v1

@@ 4.1 Model performance
Finally, we assess our model performance relative to LB23 and ZGR23 over test data in the full catalog, as a function of stellar labels. Importantly, here we do not apply the reliability cut of ZGR23 to our model error distributions, but rather only to ZGR23. Since our model is fully independent of stellar labels, we would expect to only observe error trends coming from training object density, as opposed to genuine trends coming from the labels themselves, which is indeed the case. Our model errors increase towards the wings of the stellar label distributions (top row) as the number of training objects decreases. First, our XP reconstruction errors in coefficient space are typically 0.5-1 orders of magnitude smaller than the stellar label dependent LB23 model across stellar label space, until training object density decays at the extremities. Conversely, the stellar label independent LB23 model outperforms our model across almost all of stellar label space. Second, we observe that our model is particularly more robust when applied to cool stars and low surface gravity stars. The ZGR23 error distributions below K and below dex are effectively meaningless because almost all stars in these regimes are removed by their quality cut. This can be observed by comparing the ZGR23 stellar label distributions (orange) to the XP/APOGEE cross-match stellar label distributions (green). We emphasize that this is an important advantage of our stellar label independent approach: not relying on stellar labels to simulate spectra allows our model to extend into regions of stellar parameter space where labels are unreliable. As such, with an appropriate training sample our model has the potential accurately reproduce both M-dwarf (low temperature) and giant (low surface gravity) BP/RP spectra.
@@
To summarize, our stellar label independent model outperforms the stellar label dependent transformer-based model of LB23, but underperforms in comparison to the stellar label independent LB23 model due to our comparatively harsh data compression. We also compare our model to an ‘expert’ stellar label dependent model: the deep stellar label model of ZGR23. Relative to ZGR23, we find that our stellar label independent model has specific advantages. Namely, simulating XP spectra from approximately 550-990 nm, and simulating cool stars below K and low surface gravity stars below dex.
@@
The major novelty of our sVAE model is its independence from stellar labels, opting instead to generate an XP spectrum from a latent space. This novelty is a strength, but also a potential weakness. One the one hand, in Section 4, we demonstrated that our stellar label independent model demonstrates increased performance relative to stellar label dependent models. On the other hand, stellar label dependent models are inherently useful because stellar labels are grounded in astrophysical understanding. Two questions then naturally arise: ‘What astrophysical information has the latent space learned?’ and ‘Is the latent space representation of XP spectra useful?’. We will set aside the latter question and return to it in Sections 5 and 6, focusing for now on astrophysical interpretation of the latent space.
@@ 4.2.1 Kiel tracks
With Eqs. (11) and (12), we bin and subsequently average XP spectra according to their APOGEE labels along the MS and GB relations in Kiel space to create tracks, shown in the left panel of Figure 10. We then project these spectra into our sVAE latent space, and average again to produce analogous tracks–not as a function of stellar labels, but as a function of latent variables, shown in the right panel of Figure 10. Here, we only present latent tracks in two of the most Kiel informative latent dimensions (for tracks across the full latent space, see Figure 14 in Appendix A). We observe that and function as pseudo-Kiel space, with both the MS and GB tracks exhibiting well-behaved, ordered trajectories through the latent space. We conclude that our sVAE learns the relation, despite the fact that the model has never ‘seen’ these stellar labels.
@@ 4.2.2 Metallicity tracks
In summary, our trained sVAE learns information about Gaia XP spectra which is effectively equivalent to stellar labels: and without ever having ‘seen’ any of these labels during training. However, the latent space blends its understanding of stellar labels across multiple latent dimensions. The information our sVAE learns about Gaia XP spectra contains stellar label information, but can additionally learn information from stellar spectra which is not captured by stellar labels. Furthermore, our latent space can be used to probe the actual astrophysical information content of the Gaia XP spectra, without the ambiguity produced by stellar label correlations; an important consideration for -abundance.
@@
Initially, Gavel et al. (2021) used the ExtraTrees algorithm and argued that XP spectra -abundance measurements for stars which only vary in [/M] are dangerous because XP -abundance estimates strongly depend on correlations between [/M] and other stellar properties, particularly [M/H]. This is potentially because the Gaia XP spectra are so extremely low-resolution that a change in [/M] is equivalent to a change in [M/H] due to multiple spectral lines being blurred together. Then, Witten et al. (2022) supported these findings by demonstrating that synthetic XP spectra do not contain significant -information. It is important to mention that both of these works emphasize that these difficulties are especially relevant for warm stars with 5000 K Additionally, the work of Witten et al. (2022) involved synthetic XP spectra with
More recently, Li et al. (2023) developed AspGap: a regression model which they argue is capable of producing precise [/M] estimates for red giant branch stars with K, which extends into the warm regime where -abundance estimates have been identified as problematic. Then, Hattori (2024) trained a tree-based machine learning model to estimate both [M/H] and [/M]. Additionally, Hattori (2024) proposed specific lines which contribute to -information: Na D lines (589 nm) and the Mg I line (516 nm).
@@
To contribute to answering the question of -information in Gaia XP, we separate high- and low- sequence members of the pristine labels catalog555The pristine labels catalog is strictly composed of giants with K. As such, we are probing the cool star regime where -abundance estimates are less problematic. in [M/H]-[/M] space (left panel of Figure 12) and project them into our latent space (right panel of Figure 12). To do so, we implement the -bimodality split of Patil et al. (2023), who used copulas and elicitable maps to cleanly separate the high- (purple) and low- (orange) sequences666Patil et al. (2023) define their split in [Fe/H]-[Mg/Fe] space. As such, we do the same, and subsequently present the split in [M/H]-[/M] space.. We then produce binned tracks along both sequences in [M/H]-[/M] space and project them into our latent space, similarly to the tracks in Section 4. We observe that our and latent variables function as pseudo-[M/H]-[/M] space, in which the high- and low- sequence tracks are clearly separated. However, these tracks alone are not definitive evidence of -information, because they are functions of metallicity! Therefore, we produce a third track at fixed metallicity ([M/H] , red), which only varies with [/M]. Crucially, the fixed metallicity track transitioning between both sequences in [M/H]-[/M] space is reproduced in our latent space. The high- and low- sequence tracks, with the addition of the fixed metallicity track, is compelling evidence that the Gaia XP spectra contain meaningful -information (for low-temperature stars).
To be even more pessimistic, one can also worry about and correlations with [/M] which could masquerade as [/M] trends in our latent space. Furthermore, it is also possible that the underlying extinction distributions for the high- and low- populations differ significantly. Therefore, it is important to demonstrate that our model actually learns the -bimodality without relying on stellar label and extinction relationships beyond metallicity. As such, we train four Random Forest Classifiers (sklearn.ensemble.RandomForestClassifer; Pedregosa et al., 2011) on 90% of the pristine catalog to classify the high- and low- sequences:
@@
Label classifier: trained on all stellar labels expect [/M]: and [M/H].
@@
In summary, the striking -bimodality tracks in our latent space, in combination with the significant improvement in -bimodality classification with our latent classifier relative to the [M/H], label and label + extinction classifiers, is stellar label independent evidence that the Gaia XP spectra can and should be used to estimate [/M] with K.
@@ 6 Summary, discussion and outlook
[/M] information: We provided strong evidence that the Gaia XP spectra contain meaningful -abundance information without relying on stellar label correlations, for stars with K. As such, we encourage the astronomy community to make use of the Gaia XP spectra for estimation of [/M] in this temperature regime777The extension of [/M] estimates to hotter stars in the Gaia XP spectra was not addressed in this work..
@@
Gaia XP coverage: The first implementation of our stellar label independent model was trained on a small minority of XP spectra relative to the total number of XP spectra available (). As such, we caution the application of our model to stellar populations beyond the XP/APOGEE cross-match, such as white dwarfs. This does not mean that our model cannot eventually accommodate these populations. Future implementations of our stellar label model can be trained on any/all XP spectra because our model does not require stellar labels to train on.
@@
Outlier detection: A particularly promising application of our stellar label independent model is the detection of rare sub-populations in the Gaia XP spectra via the latent space. Recently, Lucey et al. (2022) released a catalog of carbon-enhanced metal poor stars candidates (CEMPs). A preliminary application of our model to their CEMP catalog has uncovered a population of outliers. We suspect that a fraction of these outliers could be binary systems containing CEMP stars, since a major formation channel for CEMPs is binary mass transfer (e.g. Goswami et al., 2021). We will soon present a detailed analysis of this outlier population in Laroche et al. 2024 (in prep.).
In conclusion, the Gaia DR3 XP spectra is by far the largest spectroscopic survey to date. Stellar label dependent models are incapable of exploiting the entirety of the astrophysical information in the XP dataset, because they are limited by the availability of stellar labels to train on. Novel data-driven techniques must be developed to tackle this big data problem. Furthermore, as the sizes of future large-scale surveys grow, the stellar labels gap for stellar label dependent models will only widen as the relative stellar label coverage decreases (e.g. SPHEREx: Unwin et al., 2016; Roman: Zellem et al., 2022; Vera Rubin: Ivezić et al., 2019). The big-data era in astronomy is both a blessing, for discovery, and a curse, for data analysis. Our stellar label independent model is an important step towards fully data-driven modeling which can confront the intimidating amount of data which present and future large-scale surveys provide.
@@ Acknowledgments
AL and JSS would like to thank Maria Drout, Jo Bovy, Henry Leung, Gregory M. Green, Jiadong Li and Jeff Shen for valuable feedback on or help with our work. AL acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and is partially funded through a NSERC Canada Graduate Scholarship — Master’s. JSS acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) funding reference #RGPIN-2023-04849. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto.

### v2

@@ 4.1 Model performance
Finally, in Figure 9 we assess our model performance relative to LB23 and ZGR23 over test data in the full catalog, as a function of stellar labels. Importantly, here we do not apply the reliability cut of ZGR23 to our model error distributions, but rather only to ZGR23. Since our model is fully independent of stellar labels, we would expect to only observe error trends coming from training object density (panels (i)-(iii)), as opposed to genuine trends coming from the labels themselves, which is indeed the case. Our model errors increase towards the wings of the stellar label distributions (panels (iv)-(ix)) as the number of training objects decreases. First, our XP reconstruction errors in coefficient space are typically 0.5-1 orders of magnitude smaller than the stellar label dependent LB23 model across stellar label space, until training object density decays at the extremities. Conversely, the stellar label independent LB23 model outperforms our model across much of stellar label space. Second, we observe that our model is particularly more robust when applied to cool stars (panels (iv) and (vii)) and low surface gravity stars (panels (v) and (viii)). The ZGR23 error distributions below K and below dex are effectively meaningless because almost all stars in these regimes are removed by their quality cut. This can be observed by comparing the ZGR23 stellar label distributions (orange) to the XP/APOGEE cross-match stellar label distributions (green). We emphasize that this is an important advantage of our stellar label independent approach: not relying on stellar labels to simulate spectra allows our model to extend into regions of stellar parameter space where labels are unreliable. As such, with an appropriate training sample our model has the potential accurately reproduce both M-dwarf (low temperature) and giant (low surface gravity) BP/RP spectra.
@@
To summarize, our stellar label independent model outperforms the stellar label dependent transformer-based model of LB23, but underperforms in comparison to the stellar label independent LB23 model due to our comparatively harsh data compression. We also compare our model to an āexpertā stellar label dependent model: the deep stellar label model of ZGR23. Relative to ZGR23, we find that our stellar label independent model has specific advantages. Namely, simulating XP spectra from approximately 550-990 nm, and simulating cool stars below K and low surface gravity stars below dex.
@@
The major novelty of our sVAE model is its independence from stellar labels, opting instead to generate an XP spectrum from a latent space. This novelty is a strength, but also a potential weakness. One the one hand, in Section 4, we demonstrated that our stellar label independent model demonstrates increased performance relative to stellar label dependent models. On the other hand, stellar label dependent models are inherently useful because stellar labels are grounded in astrophysical understanding. Two questions then naturally arise: āWhat astrophysical information has the latent space learned?ā and āIs the latent space representation of XP spectra useful?ā. We will set aside the latter question and return to it in Sections 5 and 6, focusing for now on astrophysical interpretation of the latent space.
@@ 4.2.1 Kiel tracks
With Eqs. (11) and (12), we bin and subsequently average XP spectra according to their APOGEE labels along the MS and GB relations in Kiel space to create tracks, shown in the left panel of Figure 10. We then project these spectra into our sVAE latent space, and average again to produce analogous tracksānot as a function of stellar labels, but as a function of latent variables, shown in the right panel of Figure 10. Here, we only present latent tracks in two of the most Kiel informative latent dimensions (for tracks across the full latent space, see Figure 15 in Appendix A). We observe that and function as pseudo-Kiel space, with both the MS and GB tracks exhibiting well-behaved, ordered trajectories through the latent space. That being said, it is apparent that the GB track is slightly more stochastic than the MS track. We speculate that this is could arise from dust extinction which impacts the optical XP spectra, but not the APOGEE stellar labels derived from near-infrared spectroscopy. We conclude that our sVAE learns the relation, despite the fact that the model has never āseenā these stellar labels.
Additionally, in Figure 11 we present model reconstructed XP spectra in wavelength space for stars randomly drawn from within each bin used to define the MS and GB tracks presented in Figure 10. The percent difference for our reconstructions relative to observations are predominantly at the percent level, apart from the bluest wavelengths where reconstructions deteriorate somewhat (for an in-depth discussion on model reconstruction errors, see Section 4.1).
@@ 4.2.2 Metallicity tracks
In summary, our trained sVAE learns information about Gaia XP spectra which is effectively equivalent to stellar labels: and without ever having āseenā any of these labels during training. However, the latent space blends its understanding of stellar labels across multiple latent dimensions. The information our sVAE learns about Gaia XP spectra contains stellar label information, but can additionally learn information from stellar spectra which is not captured by stellar labels. Furthermore, our latent space can be used to probe the actual astrophysical information content of the Gaia XP spectra, without the ambiguity produced by stellar label correlations; an important consideration for -abundance.
@@
Initially, Gavel etĀ al. (2021) used the ExtraTrees algorithm and argued that XP spectra -abundance measurements for stars which only vary in [/M] are dangerous because XP -abundance estimates strongly depend on correlations between [/M] and other stellar properties, particularly [M/H]. This is potentially because the Gaia XP spectra are so extremely low-resolution that a change in [/M] is equivalent to a change in [M/H] due to multiple spectral lines being blurred together. Then, Witten etĀ al. (2022) supported these findings by demonstrating that synthetic XP spectra do not contain significant -information. It is important to mention that both of these works emphasize that these difficulties are especially relevant for warm stars with 5000 K Additionally, the work of Witten etĀ al. (2022) involved synthetic XP spectra with
More recently, Li etĀ al. (2023) developed AspGap: a regression model which they argue is capable of producing precise [/M] estimates for red giant branch stars with K, which extends into the warm regime where -abundance estimates have been identified as problematic. Then, Hattori (2024) trained a tree-based machine learning model to estimate both [M/H] and [/M]. Additionally, Hattori (2024) proposed specific lines which contribute to -information: Na D lines (589 nm) and the Mg I line (516 nm).
@@
To contribute to answering the question of -information in Gaia XP, we separate high- and low- sequence members of the pristine labels catalog666The pristine labels catalog is strictly composed of giants with K. As such, we are probing the cool star regime where -abundance estimates are less problematic. in [M/H]-[/M] space (left panel of Figure 13) and project them into our latent space (middle panel of Figure 13). To do so, we implement the -bimodality split of Patil etĀ al. (2023), who used copulas and elicitable maps to cleanly separate the high- (purple) and low- (orange) sequences777Patil etĀ al. (2023) define their split in [Fe/H]-[Mg/Fe] space. As such, we do the same, and subsequently present the split in [M/H]-[/M] space.. We then produce binned tracks along both sequences in [M/H]-[/M] space and project them into our latent space, similarly to the tracks in Section 4. We observe that our and latent variables function as pseudo-[M/H]-[/M] space, in which the high- and low- sequence tracks are clearly separated. However, these tracks alone are not definitive evidence of -information, because they are functions of metallicity! Therefore, we produce a third track at fixed metallicity ([M/H] , red), which only varies with [/M]. Crucially, the fixed metallicity track transitioning between both sequences in [M/H]-[/M] space is reproduced in our latent space. The high- and low- sequence tracks, with the addition of the fixed metallicity track, is compelling evidence that the Gaia XP spectra contain meaningful -information (for low-temperature stars).
To reinforce our argument that the latent space has learned important chemical information, we also present a sample of Gaia-Enceladus stars (e.g. Helmi etĀ al., 2018) in both [M/H]-[/M] space (left panel of Figure 13) and latent space (right panel of Figure 13). To select Gaia-Enceladus members, we implement the action diamond cuts recommended by Lane etĀ al. (2022): and for stars in the full catalog. The evident separation of the Gaia-Enceladus population from both the high- and low- sequences in the latent space (particularly across ) demonstrates that our model is sensitive to the anomalous abundances characteristic of Gaia-Enceladus.
To be even more pessimistic, one can also worry about and correlations with [/M] which could masquerade as [/M] trends in our latent space. Furthermore, it is also possible that the underlying extinction distributions for the high- and low- populations differ significantly. Therefore, it is important to demonstrate that our model actually learns the -bimodality without relying on stellar label and extinction relationships beyond metallicity. Lastly, some work has already been undertaken to present evidence for -information content in a similar manner to the above. For example, synthetic photometry from Gaia XP spectra has been used to demonstrate that [/M] appears to correlate with the Gaia colour C1M395-C1M410 (see Fig. 28 of (Gaia Collaboration etĀ al., 2022)).
As such, we train four Random Forest Classifiers (sklearn.ensemble.RandomForestClassifer; Pedregosa etĀ al., 2011) on 90% of the pristine catalog to classify the high- and low- sequences, with the intent of quantitatively demonstrating -information in a novel way. The classifiers are as follows:
@@
Label classifier: trained on all stellar labels except [/M]: and [M/H].
@@
In summary, the striking -bimodality tracks in our latent space, in combination with the significant improvement in -bimodality classification with our latent classifier relative to the [M/H], label and label + extinction classifiers, is stellar label independent evidence that the Gaia XP spectra can and should be used to estimate [/M] with K. We speculate that it may be possible to estimate [/M] above this temperature threshold if a training sample with hotter stars than APOGEE were used.
@@ 6 Summary, discussion and outlook
[/M] information: We provided strong evidence that the Gaia XP spectra contain meaningful -abundance information without relying on stellar label correlations, for stars with , while also being sensitive to the anomalous abundances of Gaia-Enceladus stars. As such, we encourage the astronomy community to make use of the Gaia XP spectra for estimation of [/M] in this temperature regime888The extension of [/M] estimates to hotter stars in the Gaia XP spectra was not addressed in this work..
@@
Gaia XP spectra coverage: The first implementation of our stellar label independent model was trained on a small minority of XP spectra relative to the total number of XP spectra available (). As such, we caution the application of our model to stellar populations beyond the XP/APOGEE cross-match, such as white dwarfs. This does not mean that our model cannot eventually accommodate these populations. Future implementations of our stellar label independent model can be trained on any/all XP spectra because our model does not require stellar labels to train on.
@@
Outlier detection: A particularly promising application of our stellar label independent model is the detection of rare sub-populations in the Gaia XP spectra via the latent space. Recently, Lucey etĀ al. (2022) released a catalog of carbon-enhanced metal poor stars candidates (CEMPs). A preliminary application of our model to their CEMP catalog has uncovered a population of outliers. We suspect that a fraction of these outliers could be binary systems containing CEMP stars, since a major formation channel for CEMPs is binary mass transfer (e.g. Goswami etĀ al., 2021). We will soon present a detailed analysis of this outlier population in Laroche et al. 2025 (in prep.).
In conclusion, the Gaia DR3 XP spectra is by far the largest spectroscopic survey to date. Stellar label dependent models are incapable of exploiting the entirety of the astrophysical information in the XP dataset, because they are limited by the availability of stellar labels to train on. Novel data-driven techniques must be developed to tackle this big data problem. Furthermore, as the sizes of future large-scale surveys grow, the stellar labels gap for stellar label dependent models will only widen as the relative stellar label coverage decreases (e.g. SPHEREx: Unwin etĀ al., 2016; Roman: Zellem etĀ al., 2022; Vera Rubin: IveziÄ etĀ al., 2019). The big-data era in astronomy is both a blessing, for discovery, and a curse, for data analysis. Our stellar label independent model is an important step towards fully data-driven modeling which can confront the intimidating amount of data which present and future large-scale surveys provide.
@@ Acknowledgments
We are thankful to the anonymous referee for providing a constructive report that helped improve the clarity and quality of our manuscript. AL and JSS would like to thank Maria Drout, Jo Bovy, Henry Leung, Gregory M. Green, Jiadong Li and Jeff Shen for valuable feedback. AL also acknowledges helpful conversations with James Lane, Nolan Koblischke and Amy Prickett which greatly improved this work. AL acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) and is (was) partially funded through a NSERC Canada Graduate Scholarship - Doctoral (Masterās). AL is also supported by the Data Sciences Institute at the University of Toronto through grant number DSI- DSFY3R1P02. JSS acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) funding reference #RGPIN-2023-04849. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Additionally, this work made use of the Python package GaiaXPy, developed and maintained by members of the Gaia DPAC, and in particular, Coordination Unit 5 (CU5), and the Data Processing Centre located at the Institute of Astronomy, Cambridge, UK (DPCI).